# Stability of discrete-symmetry flocks: sandwich state, traveling domains and motility-induced pinning

### v1

@@ III.2 Numerical simulations with droplet excitation
For $D=0.1$ and $D=0.2$, the system exhibits qualitatively different behavior compared to earlier cases: the inserted droplet forms a jammed cluster, even in the absence of external constraints on particle motion, distinct from the jamming observed in Ref. Karmakar et al. (2023). At $D=0.2$, this jammed cluster gradually dissolves as its constituent particles flip to match the spin of the background liquid, ultimately restoring the initial ordered phase (see movie07 at Ref. Chatterjee et al. (2025)). In contrast, at $D=0.1$, low motility combined with reduced noise ($\beta=1$) gives rise to the spontaneous nucleation of oppositely polarized droplets outside the original droplet, leading to the formation of multiple jammed domains. These domains appear either as horizontal clusters of $\sigma=2$ (lower side) and $\sigma=4$ (upper side), or as vertical clusters of $\sigma=1$ (left) and $\sigma=3$ (right). The extent and configuration of these jammed regions, with sharply polarized domains, are governed by the degree of self-propulsion and act as effective barriers that hinder subsequent particle motion and rearrangement (see movie08 at Ref. Chatterjee et al. (2025)).
@@ III.3 Numerical simulations in the regime of small diffusion
Figure 8: (color online) Spontaneous nucleation of polar droplets in the liquid phase. Time evolution snapshots after starting from (a) ordered ($\sigma=2$ particles, green) and (b–c) semi-ordered configurations [(b) longitudinal band of $\sigma=2$ particles, (c) transverse band of $\sigma=3$ particles] exhibit spontaneous nucleation of droplets of other spin states and subsequent destruction of the long-range ordered (LRO) phase. Parameters: $D=0.3$, $\beta=1$, $\epsilon=2.5$, (a) $\rho_{0}=5$ and (b–c) $\rho_{0}=3$. System size: (a) square domain of dimension $1024\times 1024$ and (b–c) rectangular domain of dimension $1024\times 128$. A movie (upper panel of movie09) of the same can be found at Ref. Chatterjee et al. (2025).
@@
Recent investigations of the one-species AIM Benvegnen et al. (2023); Woo and Noh (2024) and the two-species AIM with both reciprocal and non-reciprocal inter-species interactions Mangeat et al. (2025) have shown that, at low diffusion, spontaneous nucleation of droplets with opposite polarity can destabilize the flocking phase. This leads to a breakdown of global order, resulting in multiple clusters with only short-range correlations. Furthermore, it has been reported Benvegnen et al. (2023) that the nucleation time for such droplets scales as $\sim e^{D}/L^{2}$, where $L$ denotes the linear system size. In Sec. III.2, we similarly observed the spontaneous emergence of droplets within the dense liquid phase of the APM, propagating transversely to the main flow, for low diffusion and strong bias. Motivated by these findings, we now analyze the 4-state APM in the absence of any externally inserted droplet, focusing on regimes of small diffusion and high self-propulsion.
@@
Fig. 8 illustrates the spontaneous nucleation of droplets composed of particles biased to move transversely to the liquid flow, which ultimately destabilize both the polar liquid and the liquid-gas coexistence band, consistent with earlier reports for the AIM Woo and Noh (2024). Starting from a homogeneous liquid of $\sigma=2$ particles [Fig. 8(a), $t=0$], left- $(\sigma=3)$ and right-moving $(\sigma=1)$ clusters emerge spontaneously by $t=150$. These clusters grow over time $(t=300)$, alongside new spin states, and progressively destroy the long-range order (LRO) of the initial phase. At $t=3500$, the system settles into a steady state with short-range order (SRO), composed of multiple mobile domains of all four spin states. Due to the large self-propulsion bias $(\epsilon=2.5)$, these clusters align longitudinally with their direction of motion.
@@
This instability also affects the coexistence phase. Figs. 8(b–c) show that both transversely and longitudinally moving bands eventually destabilize into SRO phases, though the nucleation time differs: transverse bands break down more rapidly ($t=114$) than longitudinal ones ($t=1622$). This is expected, as the APM favors longitudinal band motion at high $\epsilon$ due to the reorientation transition Chatterjee et al. (2020); Mangeat et al. (2020). In both cases, the final SRO state is qualitatively similar at long times ($t=5532$).
@@
Unlike the AIM Woo and Noh (2024), where nucleation predominantly proceeds via counter-propagating droplets, the APM exhibits spontaneous nucleation exclusively via transversely propagating droplets, which then grow and disrupt the ordered phase. This distinction arises from the orthogonal alignment of such droplets relative to the background flow, reducing head-on collisions and enhancing persistence. At low diffusion, stochastic spin fluctuations are suppressed, allowing local spin-aligned regions to survive and coarsen through ferromagnetic alignment. The large bias $\epsilon$ further amplifies the directed motion of these droplets, enabling them to evade dominant background flow and resist dissolution. In contrast, counter-propagating droplets face stronger mechanical resistance and are more prone to decay. As discussed in Sec. III.2, even at large $\epsilon$, counter-propagating droplets tend to produce sandwich states rather than induce full phase reversal, whereas transverse droplets are more effective in destabilizing the ordered phase.
@@
We have also studied the spontaneous nucleation and breakdown of the liquid phase in a modified version of the APM, which includes a small change to the hopping mechanism inspired by the AIM variant proposed in Ref. Woo and Noh (2024). Details of this extension are provided in Appendix E.
@@
Figure 10: (color online) Motility-induced (interface) pinning in the APM. (a, c) Starting from an ordered configuration of $\sigma=2$ particles, the system forms jammed clusters in the steady state. The corresponding density plots are shown in (b) and (d). (e) Temporal evolution of the density correlation function $C_{\rho\rho}(r,t)$ (averaged over five ensembles) during the system’s transition to MIP. (f) $R(t)$ versus $t$ (on a log-log scale) showing the slow coarsening of the MIP clusters. Parameters: $D=0.3$, $\beta=2$, $\rho_{0}=5$, and $L_{x}=L_{y}=256$. A movie (movie10) of the same can be found at Ref. Chatterjee et al. (2025).
Given that one- and two-species AIM Woo and Noh (2024); Mangeat et al. (2025) display motility-induced (interface) pinning (MIP), we investigate the emergence of similar behavior in the APM under conditions of large $\beta$ (low temperature) and reduced diffusion. Initiating from an ordered configuration of $\sigma=2$ particles (green), the system evolves into a steady state featuring jammed clusters [Figs. 10(a–d)], indicating the occurrence of interface pinning at low temperature, consistent with previous findings in the AIM Woo and Noh (2024). At early stages, spontaneously nucleated droplets of opposite spin states (mainly $\sigma=1$ and $\sigma=3$) appear within the liquid region, serving as pinning centers. Consequently, the system rapidly undergoes an MIP transition, where these small droplets progressively aggregate into several narrow, dense pinned clusters over time.
The emergence of the MIP transition at small diffusion can be understood through the development of distinct timescales Woo and Noh (2024). The overall particle hopping rate scales as $4D$, and thus becomes increasingly suppressed as $D$ decreases, whereas spin-flipping processes scale as $\sim\exp(4\beta)$ and become dominant at low temperatures. Additionally, for large $\epsilon$, particle motion becomes persistent rather than diffusive. Consequently, once a jammed cluster forms, marked by the accumulation of oppositely polarized particles on either side of an interface, particles of similar polarity continue to accumulate on their respective sides due to persistent propulsion. These clusters stabilize through pinned interfaces, supported by rapid flipping and suppressed diffusion, which inhibit particle escape and strengthen the phase separation. To clearly illustrate the microscopic mechanism of pinning, we examine the behavior of a counter-propagating droplet. In Appendix F, we show how interface pinning of such a droplet can arise under conditions of low temperature and small $D$.
Figure 11: (color online) APM phase diagrams for varying diffusion. The phase diagrams are shown in: (a) the ($D$$\beta$) plane for $\rho_{0}=5$ and $\epsilon=2.5$; (b) the ($D$$\rho_{0}$) plane for $\beta=1$ and $\epsilon=2.5$; and (c) the ($D$$\epsilon$) plane for $\beta=1$ and $\rho_{0}=5$. Six distinct states are shown in the phase diagrams: gas (black diamond), long-range ordered liquid (dark-red up triangle), stripe (green circle), lane (red pentagon), short-range order (SRO, blue square), and motility-induced pinning (MIP, orange down triangle). System size: $L_{x}=L_{y}=512$.
To collate the emerging features of the APM as a function of the diffusion constant $D$ and other key parameters (average density $\rho_{0}$, particle self-propulsion velocity $\epsilon$, and inverse temperature $\beta$), we construct three phase diagrams shown in Fig. 11. For $D=1$, we recover the APM behavior reported in Ref. Chatterjee et al. (2020); Mangeat et al. (2020), where increasing either the density or $\beta$ drives the system from a coexistence phase (longitudinal lane) to a polar liquid phase exhibiting long-range order (LRO). In Fig. 11(a), we show the $D$$\beta$ phase diagram for fixed $\rho_{0}=5$ and $\epsilon=2.5$. At sufficiently large $\beta$, MIP emerges even for $D=1$, similar to the behavior observed in AIM Woo and Noh (2024), and the MIP region expands further as $D$ decreases. Lower $D$ facilitates the spontaneous formation of transverse-polarity droplets, which act as pinning centers for MIP. It is known that the nucleation time increases with increasing $D$ Benvegnen et al. (2023). For intermediate $\beta$, the spontaneous nucleation of droplets gives rise to SRO steady states, whereas for small $\beta$, as the separation between hopping and flipping timescales diminishes, the SRO state is no longer observed. Instead, depending on the temperature $(\beta^{-1})$, the system either enters a stripe state or remains in a gaseous state. The stripe state represents a liquid-gas coexistence regime and typically forms when quenching below the liquid spinodal from an ordered state, resulting in several liquid domains. This state predominantly appears at large $\epsilon$, where the strong bias suppresses transverse motion, thereby inhibiting domain coalescence (see lower panel of movie09 in Ref. Chatterjee et al. (2025)).
The $D$$\rho_{0}$ phase diagram [Fig. 11(b)] is shown for $\beta=1$ and $\epsilon=2.5$. This diagram is largely dominated by the short-range order (SRO) phase, with the SRO region expanding and the liquid region diminishing as $D$ is decreased down to $D=0.2$, since lower $D$ enhances the spontaneous emergence of transverse-polarity droplets. At very small diffusion ($D=0.1$), we observe a transition from a gas phase at low densities to a MIP phase at higher densities. A similar occurrence of a MIP-like jammed phase for $D=0.1$ is discussed in Sec. III.2, in the context of Fig. 7. This onset of MIP at small $D$ is driven by two competing timescales: a moderately fast flipping process for $\beta=1$ and a slow hopping process associated with $D=0.1$.
The $D$$\epsilon$ phase diagram [Fig. 11(c)] is plotted for $\beta=1$, $\rho_{0}=5$, and relatively large values of $\epsilon$, as the SRO phase is absent at small $\epsilon$. Although droplet nucleation still occurs at small $\epsilon$, the resulting droplets cannot persist due to enhanced hopping in directions orthogonal to the bias. Moreover, the critical droplet size needed to destabilize the liquid increases with decreasing $\epsilon$ (see Sec. III.2). At $D=1$, increasing $\epsilon$ causes a transition from the liquid phase to a stripe state. Since hopping along the biased directions becomes dominant at high velocities, the transverse extent of the liquid region becomes compressed. The SRO phase emerges in the range $0.2\leq D\leq 0.7$ across most values of $\epsilon$, except for $D=0.1$, where the MIP state becomes dominant.
@@
Taken together, the phase diagram of the APM Chatterjee et al. (2020); Mangeat et al. (2020) becomes substantially richer when the diffusion constant $D$ is varied. Beyond the conventional APM phases, we identify two novel states: a short-range ordered (SRO) phase and a motility-induced pinning (MIP) phase, along with a stripe coexistence region not previously reported. Including these with the gas, liquid, and liquid-gas coexistence phases yields a comprehensive understanding of the complex phase behavior in the APM.
@@ IV Discussion
Motivated by recent studies highlighting the fragility of the polar ordered phase in various flocking models Besse et al. (2022); Benvegnen et al. (2023); Woo and Noh (2024); Karmakar et al. (2024); Mangeat et al. (2025), we have investigated the stability of ordered phases in a discrete-symmetry flocking model, specifically the 4-state APM Chatterjee et al. (2020); Mangeat et al. (2020). Our analysis began with an examination of how small counter-propagating and transversely propagating droplets perturb the polar liquid phase. Interactions between these droplets and the background polar flocks generate local disturbances that can result in either a full reversal of the flock or the emergence of a distinct intermediate configuration termed as the sandwich state, where a reversed domain becomes embedded between two domains of the original flock. The balance between these two outcomes is found to depend critically on the self-propulsion speed $(\epsilon)$, the inverse temperature $(\beta)$, average particle density $(\rho_{0})$, and the underlying lattice anisotropy. At high $\beta$, strong spin alignment and reduced transverse fluctuations favor the formation of sandwich states. Conversely, at lower $\beta$, complete reversals become likely within a finite intermediate range of $\epsilon$. An increase in $\epsilon$ reduces transverse diffusion, facilitating the formation of sandwich states by counter-propagating droplets and stabilizing transverse droplets by minimizing leakage across the interface. Additionally, higher $\epsilon$ lowers the threshold droplet size required to initiate either process.
@@
We further confirm that the APM, similar to the AIM Woo and Noh (2024); Mangeat et al. (2025), fails to support long-range polar order over a wide region of its phase space. Under conditions of low diffusion and high directional bias, the APM exhibits spontaneous nucleation of transversely propagating droplets, which fragment a globally ordered flock into multiple smaller domains corresponding to the four internal spin states. At very low diffusion and large $\beta$, this process gives rise to a motility-induced pinning transition, leading to jammed configurations characterized by pinned interfaces, even in the absence of quenched disorder. This phenomenon is consistent with behavior observed in both single-species and multi-species AIM systems Woo and Noh (2024); Mangeat et al. (2025). System size also plays a crucial role in determining metastability. For counter-propagating droplets, increasing the transverse system size suppresses full reversals and favors the emergence of sandwich states. On the other hand, larger systems facilitate faster spontaneous droplet nucleation, particularly for transversely moving small droplets Benvegnen et al. (2023).
@@
Altogether, these results reveal a rich interplay between temperature, diffusion, self-propulsion velocity, and system size in shaping the stability of phases in discrete flocking models such as the APM. We hope that our findings, in conjunction with previous works Besse et al. (2022); Benvegnen et al. (2023); Woo and Noh (2024); Karmakar et al. (2024); Mangeat et al. (2025), encourage further exploration into the stability of polar order in a broad range of active matter systems.

### v2

@@ III.2 Numerical simulations with droplet excitation
For $D=0.1$ and $D=0.2$, the system exhibits qualitatively different behavior compared to earlier cases: the inserted droplet forms a jammed cluster, even in the absence of external constraints on particle motion, distinct from the jamming observed in Ref. karmakar2023jamming. At $D=0.2$, this jammed cluster gradually dissolves as its constituent particles flip to match the spin of the background liquid, ultimately restoring the initial ordered phase (see movie07 at Ref. zenodo). In contrast, at $D=0.1$, low motility combined with reduced noise ($\beta=1$) gives rise to the spontaneous nucleation of oppositely polarized droplets outside the original droplet, leading to the formation of multiple jammed domains. These domains appear either as horizontal clusters of $\sigma=2$ (lower side) and $\sigma=4$ (upper side), or as vertical clusters of $\sigma=1$ (left) and $\sigma=3$ (right). The extent and configuration of these jammed regions, with sharply polarized domains, are governed by the degree of self-propulsion and act as effective barriers that hinder subsequent particle motion and rearrangement (see movie08 at Ref. zenodo).
@@ III.3 Numerical simulations in the regime of small diffusion
Figure 8: (color online) Spontaneous nucleation of polar droplets in the liquid phase. Time evolution snapshots after starting from (a) ordered ($\sigma=2$ particles, green) and (b–c) semi-ordered configurations [(b) longitudinal band of $\sigma=2$ particles, (c) transverse band of $\sigma=3$ particles] exhibit spontaneous nucleation of droplets of other spin states and subsequent destruction of the long-range ordered (LRO) phase. Parameters: $D=0.3$, $\beta=1$, $\epsilon=2.5$, (a) $\rho_{0}=5$ and (b–c) $\rho_{0}=3$. System size: (a) square domain of dimension $1024\times 1024$ and (b–c) rectangular domain of dimension $1024\times 128$. A movie (upper panel of movie09) of the same can be found at Ref. zenodo.
@@
Recent investigations of the one-species AIM benvegnen2023meta; jd2024MIP and the two-species AIM with both reciprocal and non-reciprocal inter-species interactions TSAIM have shown that, at low diffusion, spontaneous nucleation of droplets with opposite polarity can destabilize the flocking phase. This leads to a breakdown of global order, resulting in multiple clusters with only short-range correlations. Furthermore, it has been reported benvegnen2023meta that the nucleation time for such droplets scales as $\sim e^{D}/L^{2}$, where $L$ denotes the linear system size. In Sec. III.2, we similarly observed the spontaneous emergence of droplets within the dense liquid phase of the APM, propagating transversely to the main flow, for low diffusion and strong bias. Motivated by these findings, we now analyze the 4-state APM in the absence of any externally inserted droplet, focusing on regimes of small diffusion and high self-propulsion.
@@
Fig. 8 illustrates the spontaneous nucleation of droplets composed of particles biased to move transversely to the liquid flow, which ultimately destabilize both the polar liquid and the liquid-gas coexistence band, consistent with earlier reports for the AIM jd2024MIP. Starting from a homogeneous liquid of $\sigma=2$ particles [Fig. 8(a), $t=0$], left- $(\sigma=3)$ and right-moving $(\sigma=1)$ clusters emerge spontaneously by $t=150$. These clusters grow over time $(t=300)$, alongside new spin states, and progressively destroy the long-range order (LRO) of the initial phase. At $t=3500$, the system settles into a steady state with short-range order (SRO), composed of multiple mobile domains of all four spin states. Due to the large self-propulsion bias $(\epsilon=2.5)$, these clusters align longitudinally with their direction of motion.
@@
This instability also affects the coexistence phase. Figs. 8(b–c) show that both transversely and longitudinally moving bands eventually destabilize into SRO phases, though the nucleation time differs: transverse bands break down more rapidly ($t=114$) than longitudinal ones ($t=1622$). This is expected, as the APM favors longitudinal band motion at high $\epsilon$ due to the reorientation transition chatterjee2020flocking; mangeat2020flocking. In both cases, the final SRO state is qualitatively similar at long times ($t=5532$).
@@
Unlike the AIM jd2024MIP, where nucleation predominantly proceeds via counter-propagating droplets, the APM exhibits spontaneous nucleation exclusively via transversely propagating droplets, which then grow and disrupt the ordered phase. This distinction arises from the orthogonal alignment of such droplets relative to the background flow, reducing head-on collisions and enhancing persistence. At low diffusion, stochastic spin fluctuations are suppressed, allowing local spin-aligned regions to survive and coarsen through ferromagnetic alignment. The large bias $\epsilon$ further amplifies the directed motion of these droplets, enabling them to evade dominant background flow and resist dissolution. In contrast, counter-propagating droplets face stronger mechanical resistance and are more prone to decay. As discussed in Sec. III.2, even at large $\epsilon$, counter-propagating droplets tend to produce sandwich states rather than induce full phase reversal, whereas transverse droplets are more effective in destabilizing the ordered phase.
@@
We have also studied the spontaneous nucleation and breakdown of the liquid phase in a modified version of the APM, which includes a small change to the hopping mechanism inspired by the AIM variant proposed in Ref. jd2024MIP. Details of this extension are provided in Appendix F.
@@
We now examine the SRO steady state that emerges due to spontaneous droplet nucleation in the APM liquid phase and estimate the typical cluster sizes as shown in Fig. 9. Figs. 9(a–b) respectively display an exponential decay in the magnetization and density correlation functions over time, suggesting that the polar flock transitions into a new steady state with spatially disordered local domains, indicating short-range order resulting from droplet nucleation. Fig. 9(a) illustrates the time evolution of the two-point equal-time magnetization correlation function, $C_{mm}(r,t)=\sum_{i}m_{i}^{\sigma=2}(t)m_{i+r}^{\sigma=2}(t)/\rho_{0}^{2}L_{x}L_{y}$, where initially $C_{mm}(r)\sim\text{const.}$ at $t=0$, and gradually transitions to $C_{mm}(r)\sim\exp(-r/\xi)$ as time increases, with $\xi$ representing the correlation length. The typical size of the short-range clusters, $\xi=51$, is estimated from the exponential decay of the density correlation function, $C_{\rho\rho}(r)=\sum_{i}\langle\rho_{i}\rho_{i+r}\rangle/\rho_{0}^{2}L_{x}L_{y}$, with spatial separation $r$, where $\langle\dots\rangle$ indicates a steady-state temporal average [Fig. 9(b)]. This value of $\xi$ is found to be independent of the system size [Fig. 9(c)], consistent with observations in the AIM model jd2024MIP.
Figure 10: (color online) Motility-induced (interface) pinning in the APM. (a, c) Starting from an ordered configuration of $\sigma=2$ particles, the system forms jammed clusters in the steady state. The corresponding density plots are shown in (b) and (d). (e) Temporal evolution of the density correlation function $C_{\rho\rho}(r,t)$ (averaged over five ensembles) during the system’s transition to MIP. (f) $R(t)$ versus $t$ (on a log-log scale) showing the slow coarsening of the MIP clusters. Parameters: $D=0.3$, $\beta=2$, $\rho_{0}=5$, and $L_{x}=L_{y}=256$. A movie (movie10) of the same can be found at Ref. zenodo.
Given that one- and two-species AIM jd2024MIP; TSAIM display motility-induced (interface) pinning (MIP), we investigate the emergence of similar behavior in the APM under conditions of large $\beta$ (low temperature) and reduced diffusion. Initiating from an ordered configuration of $\sigma=2$ particles (green), the system evolves into a steady state featuring jammed clusters [Figs. 10(a–d)], indicating the occurrence of interface pinning at low temperature, consistent with previous findings in the AIM jd2024MIP. At early stages, spontaneously nucleated droplets of opposite spin states (mainly $\sigma=1$ and $\sigma=3$) appear within the liquid region, serving as pinning centers. Consequently, the system rapidly undergoes an MIP transition, where these small droplets progressively aggregate into several narrow, dense pinned clusters over time.
The emergence of the MIP transition at small diffusion can be understood through the development of distinct timescales jd2024MIP. The overall particle hopping rate scales as $4D$, and thus becomes increasingly suppressed as $D$ decreases, whereas spin-flipping processes scale as $\sim\exp(4\beta)$ and become dominant at low temperatures. Additionally, for large $\epsilon$, particle motion becomes persistent rather than diffusive. Consequently, once a jammed cluster forms, marked by the accumulation of oppositely polarized particles on either side of an interface, particles of similar polarity continue to accumulate on their respective sides due to persistent propulsion. These clusters stabilize through pinned interfaces, supported by rapid flipping and suppressed diffusion, which inhibit particle escape and strengthen the phase separation. To clearly illustrate the microscopic mechanism of pinning, we examine the behavior of a counter-propagating droplet. In Appendix G, we show how interface pinning of such a droplet can arise under conditions of low temperature and small $D$.
@@
[…]
@@
Taken together, the phase diagram of the APM chatterjee2020flocking; mangeat2020flocking becomes substantially richer when the diffusion constant $D$ is varied. Beyond the conventional APM phases, we identify two novel states: a short-range ordered (SRO) phase and a motility-induced pinning (MIP) phase, along with a stripe coexistence region not previously reported. Including these with the gas, liquid, and liquid-gas coexistence phases yields a comprehensive understanding of the complex phase behavior in the APM.
@@ IV Discussion
Motivated by recent studies highlighting the fragility of the polar ordered phase in various flocking models besse2022metastability; benvegnen2023meta; jd2024MIP; karmakar2024consequence; TSAIM, we have investigated the stability of ordered phases in a discrete-symmetry flocking model, specifically the 4-state APM chatterjee2020flocking; mangeat2020flocking. Our analysis began with an examination of how small counter-propagating and transversely propagating droplets perturb the polar liquid phase. Interactions between these droplets and the background polar flocks generate local disturbances that can result in either a full reversal of the flock or the emergence of a distinct intermediate configuration termed as the sandwich state, where a reversed domain becomes embedded between two domains of the original flock. The balance between these two outcomes is found to depend critically on the self-propulsion speed $(\epsilon)$, the inverse temperature $(\beta)$, average particle density $(\rho_{0})$, and the underlying lattice anisotropy. At high $\beta$, strong spin alignment and reduced transverse fluctuations favor the formation of sandwich states. Conversely, at lower $\beta$, complete reversals become likely within a finite intermediate range of $\epsilon$. An increase in $\epsilon$ reduces transverse diffusion, facilitating the formation of sandwich states by counter-propagating droplets and stabilizing transverse droplets by minimizing leakage across the interface. Additionally, higher $\epsilon$ lowers the threshold droplet size required to initiate either process.
@@
We further confirm that the APM, similar to the AIM jd2024MIP; TSAIM, fails to support long-range polar order over a wide region of its phase space. Under conditions of low diffusion and high directional bias, the APM exhibits spontaneous nucleation of transversely propagating droplets, which fragment a globally ordered flock into multiple smaller domains corresponding to the four internal spin states. At very low diffusion and large $\beta$, this process gives rise to a motility-induced pinning transition, leading to jammed configurations characterized by pinned interfaces, even in the absence of quenched disorder. This phenomenon is consistent with behavior observed in both single-species and multi-species AIM systems jd2024MIP; TSAIM. System size also plays a crucial role in determining metastability. For counter-propagating droplets, increasing the transverse system size suppresses full reversals and favors the emergence of sandwich states. On the other hand, larger systems facilitate faster spontaneous droplet nucleation, particularly for transversely moving small droplets benvegnen2023meta.
@@
Altogether, these results reveal a rich interplay between temperature, diffusion, self-propulsion velocity, and system size in shaping the stability of phases in discrete flocking models such as the APM. We hope that our findings, in conjunction with previous works besse2022metastability; benvegnen2023meta; jd2024MIP; karmakar2024consequence; TSAIM, encourage further exploration into the stability of polar order in a broad range of active matter systems. Beyond theory, our observed phases should be testable in dry active systems where motion is constrained to a few discrete directions by design or confinement. Examples include Quincke rollers guided by square electrode grids, vibrated polar grains on grooved square substrates, NESW-programmed robot swarms, and bidirectional corridor flows (such as ants or pedestrians) geyer2019freezing; bricard2013emergence; lei2023exploring; poissonnier2019experimental; feliciani2016empirical.